# SVL: Spike-based Vision-language Pretraining for Efficient 3D Open-world Understanding

### v1

@@ Abstract
Spiking Neural Networks (SNNs) provide an energy-efficient way to extract 3D spatio-temporal features. However, existing SNNs still exhibit a significant performance gap compared to Artificial Neural Networks (ANNs) due to inadequate pre-training strategies. These limitations manifest as restricted generalization ability, task specificity, and a lack of multimodal understanding, particularly in challenging tasks such as multimodal question answering and zero-shot 3D classification. To overcome these challenges, we propose a Spike-based Vision-Language (SVL) pretraining framework that empowers SNNs with open-world 3D understanding while maintaining spike-driven efficiency. SVL introduces two key components: (i) Multi-scale Triple Alignment (MTA) for label-free triplet-based contrastive learning across 3D, image, and text modalities, and (ii) Re-parameterizable Vision-Language Integration (Rep-VLI) to enable lightweight inference without relying on large text encoders. Extensive experiments show that SVL achieves a top-1 accuracy of 85.4% in zero-shot 3D classification, surpassing advanced ANN models, and consistently outperforms prior SNNs on downstream tasks, including 3D classification (+6.1%), DVS action recognition (+2.1%), 3D detection (+1.1%), and 3D segmentation (+2.1%) with remarkable efficiency. Moreover, SVL enables SNNs to perform open-world 3D question answering, sometimes outperforming ANNs. To the best of our knowledge, SVL represents the first scalable, generalizable, and hardware-friendly paradigm for 3D open-world understanding, effectively bridging the gap between SNNs and ANNs in complex open-world understanding tasks. Code is available Here.
@@ 1 Introduction
Bio-inspired Spiking Neural Networks (SNNs) offer an efficient approach to learning superior representations from sparse 3D geometric data (e.g., event streams and point clouds) [1], owing to their distinctive spike-driven nature [2] and spatio-temporal processing capabilities [3]. For instance, the Speck [4] chip uses event-by-event sparse processing to handle 3D input data, with operational power consumption as low as 0.7 mW. However, existing SNNs [5; 6; 7] exhibit a significant performance gap compared to ANNs, and remain task-dependent, lacking both generalizable representations and the ability to achieve multimodal understanding in 3D open-world scenarios.
For instance, when deploying SNNs in real-world scenarios [4], they may struggle to generalize to input data from unseen categories not present in the training set. This highlights the critical need to develop robust pretraining strategies to enhance the visual representation capabilities and adaptability of SNNs. Existing methods, such as STDP-based initialization [8] and knowledge distillation in SpikeBert and SpikeCLIP [9; 10; 11], refine spike-based representations, while SpikformerV2 and Spike-driven Transformer V3 [6; 7] employ masked image modeling to improve scalability. However, these approaches [8] either lose effectiveness as dataset complexity increases, demand substantial computational resources [7], which limits their feasibility for neuromorphic hardware deployment, or lack multimodal integration [9]. Moreover, pre-trained models often exhibit inadequate visual representation capabilities and limited transferability, restricting unified applicability to downstream tasks [6].
Another challenge is the limited availability of annotated 3D datasets, as the creation of such datasets is both labor-intensive and error-prone, rendering it often impractical for large-scale, real-world applications [12]. In response, Vision-Language Models (VLMs) [13] have been employed to explore the transfer of knowledge gleaned from extensive 2D datasets to facilitate open-world 3D understanding. However, most VLMs [14; 15; 16] depend heavily on large-scale text encoders during inference, which imposes substantial limitations on the practicality of hardware deployment.
@@
In this paper, we introduce a universal Spike-based Vision-Language pretraining framework (SVL) that enhances SNNs’ capability for open-world multimodal 3D understanding while maintaining efficient spike-driven inference. As shown in Fig. 1, our framework incorporates two key innovations: (i) Multi-scale Triple Alignment (MTA), which enables label-free triplet representation learning for capturing geometric properties of 3D data, and (ii) Re-parameterizable Vision-language Integration (Rep-VLI), which facilitates efficient deployment by addressing the computational overhead of text-driven inference. Moreover, our framework uses CLIP for its strong generalization. During pre-training, we freeze CLIP and train the 3D encoder by aligning 3D features with CLIP’s textual and visual features via contrastive learning. The pre-trained 3D model can then be fine-tuned for downstream tasks. To demonstrate SVL’s efficacy, we rigorously evaluated our models on a wide range of 3D tasks, including open-world understanding and downstream applications. We also highlight SVL’s versatility through its potential in generative tasks like 3D object captioning and open-world question answering, showcasing its broad applicability in 3D vision domains. Our main contribution can be summarized as:
  • We introduce SVL, a universal spike-based vision-language pretraining framework that enables SNNs to understand open-world multimodal 3D information while maintaining efficient spike-driven inference.
  • We propose two key technical innovations: (i) Multi-scale Triple Alignment (MTA), a label-free triplet learning mechanism for capturing geometric properties of 3D data across different scales, and (ii) Re-parameterizable Vision-language Integration (Rep-VLI), which achieves lightweight inference by reducing the computational overhead of text encoder.
@@ 2 Related Works
Numerous pretraining methods are proposed for spike-based representation learning. [8] utilized spike-timing-dependent plasticity [17] to initialize SNNs, enhancing the model’s robustness and training speed. While this approach has been successful on simple datasets with shallow networks, its effectiveness diminishes as the complexity of the datasets and networks increases. To address these issues, SpikeBert and SpikeCLIP [9; 10; 11] employ a two-stage knowledge distillation process from ANNs to enhance spike-based representations for complex downstream tasks. However, these methods rely on ANN weight initialization, limiting structural flexibility. Additionally, they use LayerNorm, which hinders neuromorphic hardware deployment. SpikformerV2 and Spike-driven Transformer V3 [6; 7; 18] apply a masked image modeling approach to address the performance degradation in SNNs as the model scales up. However, they require substantial storage and computational resources, and the lack of multimodal integration, particularly language guidance, limits their effectiveness in open-world understanding tasks.
@@
aim to align image and text embeddings for cross-modal transfer, with CLIP [13] being a seminal work that uses contrastive learning for zero-shot classification. Building on this foundation, subsequent methods have expanded cross-modal alignment to include other modalities. These approaches typically fall into two categories: dual-encoder and triple-encoder frameworks. Dual-encoder fine-tune both visual and textual encoders [11; 19; 20]. Triple-encoder frameworks incorporate additional modality-specific encoders [14; 15; 21], which combine triple models to achieve open-world understanding. This architecture is highly flexible, making it suitable for a variety of downstream tasks [12; 16]. However, triple-encoder frameworks still rely on large text encoders during inference, hindering hardware deployment.
@@
is to learn better representations from sparse and irregular 3D geometric data (e.g., event and point cloud). Currently, there exist two main approaches. Point-based methods [22; 23] utilize the PointNet series to directly extract geometric features from raw 3D data and make predictions. Voxel-based methods [24] first convert the 3D data into regular voxels and then use 3D sparse convolutions for feature extraction. Nevertheless, the improved accuracy is often accompanied by increased computational costs, limiting its applicability in practical systems Hence, numerous studies [25; 26; 27] in the SNN field combine spiking neurons with point-based methods like PointNet [22] for low energy consumption edge computing. These methods are too simplistic to apply to various tasks and datasets. It was not until E-3DSNN [5] used spike sparse convolution to achieve high performance in numerous 3D tasks with spike-driven nature. Our approach is orthogonal to the aforementioned 3D encoders. Pre-training with SVL can enhance the visual extraction capabilities of the 3D encoders, thereby improving their performance in 3D recognition tasks. SVL can also endow them with the ability to achieve open-world understanding while retaining their spike-driven nature.
@@ 3 Preliminaries
are inspired by the dynamics of biological neurons [3; 28], which are the fundamental units of Spiking Neural Networks (SNNs). Among these, the Leaky Integrate-and-Fire (LIF) neuron is the most widely used due to its balance between biological plausibility and computational efficiency [3]. We begin by translating the LIF spiking neuron into an iterative expression using the Euler method [29], which is described as follows:
@@
Directly training the above LIF-based SNNs requires the use of backpropagation through time (BPTT) [29], resulting in a time complexity of $\mathcal{O}(LT)$, where $L$ and $T$ are the number of layers and time steps. This significantly increases both the training time and memory requirements. To mitigate this issue, we use the Integer LIF Spiking Neuron.
@@
is incorporated into our SVL to reduce the quantization error, training time, and memory [6; 30; 31], which allows us to rewrite Eq. (2) as:
@@
[…]
CLIP
employs a contrastive learning framework to forge associations between images and textual descriptions [13]. It leverages a vast dataset comprising 400 million image-text pairs, training an image encoder based on either a ResNet [32] or Vision Transformer [33] architecture, and a text encoder utilizing the Transformer model [34]. These encoders work together to project both images and text into a unified embedding space. This training paradigm enables CLIP to perform zero-shot classification, allowing it to identify images solely based on textual descriptions, even without category-specific training. In this study, we aim to utilize CLIP’s semantically rich feature space to support spike-based encoders, facilitating energy-efficient 3D open-world understanding with a spike-driven nature.
@@ 4 Method
In this part, we present the 3D input representation, such as point clouds and event streams. Event streams, in particular, require special handling. We define them as $E_{i}=(x_{i},y_{i},t_{i},p_{i})$. Using a sliding window technique [23; 25], we convert event streams into an event cloud, formulated as:
$E_{i}=(x_{i},y_{i},z_{i})\quad\text{where}\quad z_{i}=\frac{t_{i}-t_{\text{min}}}{t_{\text{max}}-t_{\text{min}}},$
By doing so, we treat event streams as a distinct kind of spatio-temporal point cloud. This allows us to consider both point clouds and event streams as collections of points, denoted by ${D}^{t}=\{\mathcal{P},\mathcal{I}\}$. This includes voxel sets ${D}_{k}^{t}=\{\mathcal{P}_{k}^{t},\mathcal{F}_{k}^{t}\}$, where $\mathcal{P}_{k}^{t}\in\mathbb{R}^{3}$ represents the 3D coordinates and $\mathcal{I}_{k}^{t}\in\mathbb{R}^{D}$ indicate the features across $d$ channels at the time step $t$. Following this, we utilize our I-LIF spiking neuron to encode these 3D inputs into spatio-temporal spike trains, which are then transmitted to the spike encoder.
@@ 4.2 Multi-scale Triple Alignment
To leverage the open-world recognition capabilities of the pre-trained CLIP model [13], we align the spike firing rate $\mathcal{F}^{S}/T$ with the text embeddings $\mathcal{F}^{T}$ obtained from CLIP, using a spike-text tuple $\mathcal{B}_{i}=\{{T}^{t}_{i},\mathcal{D}^{t}_{i}\}$ as input. The core idea is to bring the feature centroids of 3D instances and their corresponding text prompts closer together in the embedding space. To achieve this, we compute the InfoNCE loss [35] between the mean spike trains and the text features, as follows:
($\mathcal{L}^{\text{NCE}}_{(\textbf{S},\textbf{T})}=-\frac{1}{2|\mathcal{B}|}\sum_{i}^{|\mathcal{B}|}\log\frac{e^{\tau\mathbf{x}_{i}\cdot\mathbf{y}_{i}}}{\sum_{j}^{|\mathcal{B}|}e^{\tau\mathbf{x}_{i}\cdot\mathbf{y}_{j}}}+\log\frac{e^{\tau\mathbf{x}_{i}\cdot\mathbf{y}_{i}}}{\sum_{j}^{|\mathcal{B}|}e^{\tau\mathbf{x}_{j}\cdot y_{i}}},$ (5)
where $\mathbf{x}_{i}=\frac{\mathcal{F}^{S}/T}{||\mathcal{F}^{S}/T||2}$ and $\mathbf{y}_{i}=\frac{\mathcal{F}^{T}}{||\mathcal{F}^{T}||_{2}}$ represent the normalized spike and text features, respectively. The indices $i$ and $j$ are used for sampling, the dot product ($\cdot$) denotes cosine similarity between vectors, and $\tau$ is a learnable temperature parameter.
@@
($\mathcal{L}^{\text{NCE}}_{(\textbf{S},\textbf{I})}=-\frac{1}{2|\mathcal{C}|}\sum_{i}^{|\mathcal{C}|}\log\frac{e^{\tau\mathbf{a}_{i}\cdot\mathbf{b}_{i}}}{\sum_{j}^{|\mathcal{C}|}e^{\tau\mathbf{a}_{i}\cdot\mathbf{b}_{j}}}+\log\frac{e^{\tau\mathbf{a}_{i}\cdot\mathbf{b}_{i}}}{\sum_{j}^{|\mathcal{B}|}e^{\tau\mathbf{a}_{j}\cdot b_{i}}},$ (6)
where $\mathcal{C}_{i}=\{{I}^{t}_{i},\mathcal{D}^{t}_{i}\}$ a spike-image tuple, $\mathbf{a}_{i}=\frac{\mathcal{F}^{S}/T}{||\mathcal{F}^{S}/T||2}$ and $\mathbf{b}_{i}=\frac{\mathcal{F}^{I}}{||\mathcal{F}^{I}||_{2}}$ represent the normalized spike and text features, respectively. However, this approach resulted in overly coarse alignment granularity, failing to account for the fine-grained and tightly coupled alignment between spikes and images. To address this, we incorporate the MSE loss on the basis of the InfoNCE loss to enhance the alignment granularity. The alignment objective between spike trains and images, which is formulated as follows:
$\mathcal{L}_{(\mathbf{S},\mathbf{I})}^{\text{MSE}}=\sum_{i}^{|\mathcal{C}|}||\mathcal{F}_{i}^{\textbf{S}}-\mathcal{F}_{i}^{\textbf{I}}||^{2},$ (7)
@@
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are hyperparameters that balance the influence of image features, text features, and spike trains. In this paper, the default setting of both is 1.

### v2

# SVL: Empowering Spiking Neural Networks for Efficient 3D Open-World Understanding

@@ Abstract
Spiking Neural Networks (SNNs) offer an energy–efficient route to 3D spatio–temporal perception, yet they lag behind Artificial Neural Networks (ANNs) due to weak pretraining and heavy inference stacks, limiting generalization and multimodal reasoning (e.g., zero–shot 3D classification and open–world QA). We present a universal Spike–based Vision–Language pretraining framework (SVL) that equips SNNs with open–world 3D understanding while preserving end–to–end spike efficiency. SVL comprises two core components: (i) Multi–scale Triple Alignment (MTA), a label–free triplet contrastive objective aligning 3D, image, and text; and (ii) Re–parameterizable Vision–Language Integration (Rep–VLI), which converts offline text embeddings into lightweight weights for text–encoder–free inference. Moreover, we present the first fully spike–driven point Transformer, Spike-driven PointFormer, whose 3D spike–driven self–attention (3D-SDSA) reduces interactions to sparse additions, enabling faster, more efficient training. Extensive experiments show that SVL attains strong zero–shot 3D classification (85.4% top–1) and consistently outperforms prior SNNs on downstream tasks (e.g., +6.1% 3D cls, +2.1% DVS actions, +1.1% detection, +2.1% segmentation) while enabling open–world 3D question answering, sometimes outperforming ANNs. To the best of our knowledge, SVL represents the first scalable, generalizable, and hardware-friendly paradigm for 3D open-world understanding, effectively bridging the gap between SNNs and ANNs in complex open-world understanding tasks. Code is available at SVL.
Machine Learning, ICML, Spiking Neural Networks, 3D Open-World Understanding
@@ 1 Introduction
Bio-inspired Spiking Neural Networks (SNNs) offer an efficient approach to learning superior representations from sparse 3D geometric data (e.g., event streams and point clouds) (Roy et al., 2019), owing to their distinctive spike-driven nature (Pei et al., 2019) and spatio-temporal processing capabilities (Maass, 1997). For instance, the Speck (Yao et al., 2024) chip uses event-by-event sparse processing to handle 3D input data, with operational power consumption as low as 0.7 mW. However, existing SNNs (Qiu et al., 2025a; Yao et al., 2025; Zhou et al., 2024) exhibit a significant performance gap compared to ANNs, and remain task-dependent, lacking both generalizable representations and the ability to achieve multimodal understanding in 3D open-world scenarios.
For instance, when deploying SNNs in real-world scenarios (Yao et al., 2024), they may struggle to generalize to input data from unseen categories not present in the training set. This highlights the critical need to develop robust pretraining strategies to enhance the visual representation capabilities and adaptability of SNNs. Existing methods, such as STDP-based initialization (Lee et al., 2018) and knowledge distillation in SpikeBert and SpikeCLIP (Lv et al., 2023; Bal and Sengupta, 2024; Lv et al., 2025), refine spike-based representations, while SpikformerV2 and Spike-driven Transformer V3 (Yao et al., 2025; Zhou et al., 2024) employ masked image modeling to improve scalability. However, these approaches (Lee et al., 2018) either lose effectiveness as dataset complexity increases, demand substantial computational resources (Zhou et al., 2024), which limits their feasibility for neuromorphic hardware deployment, or lack multimodal integration (Lv et al., 2023). Moreover, pre-trained models often exhibit inadequate visual representation capabilities and limited transferability, restricting unified applicability to downstream tasks (Yao et al., 2025).
@@
Another challenge is the limited availability of annotated 3D datasets, as the creation of such datasets is both labor-intensive and error-prone, rendering it often impractical for large-scale, real-world applications (Xu et al., 2024; Liu et al., 2025). In response, Vision-Language Models (VLMs) (Radford et al., 2021; Miao et al., 2025; Tan et al., 2025) have been employed to explore the transfer of knowledge gleaned from extensive 2D datasets to facilitate open-world 3D understanding. However, most VLMs (Xue et al., 2023, 2024; Liu et al., 2023b) depend heavily on large-scale text encoders during inference, which imposes substantial limitations on the practicality of hardware deployment.
In this paper, we introduce a universal Spike-based Vision-Language pretraining framework (SVL) that enhances SNNs’ capability for open-world multimodal 3D understanding while maintaining efficient spike-driven inference. As shown in Fig. 1, SVL incorporates two key innovations: (i) Multi-scale Triple Alignment (MTA), which enables label-free triplet representation learning to capture the geometric properties of 3D data, and (ii) Re-parameterizable Vision–Language Integration (Rep–VLI), which converts offline text embeddings into lightweight weights for text–encoder–free deployment. We leverage CLIP for its strong generalization: during pretraining, CLIP is frozen and a spike-driven 3D encoder is aligned to CLIP’s image/text spaces via contrastive learning; the pre-trained 3D model is then fine-tuned for downstream tasks. In addition to SVL , we present the first fully spike–driven point Transformer, Spike-driven PointFormer, whose 3D spike–driven self–attention (3D-SDSA) reduces interactions to sparse additions, enabling faster, more efficient training that supports large–scale pretraining and generalizes broadly across 3D tasks. Our main contribution can be summarized as:
  • We propose two key innovations about SVL: (i) Multi-scale Triple Alignment (MTA), a label-free triplet learning mechanism for capturing geometric properties of 3D data across different scales, and (ii) Re-parameterizable Vision-language Integration (Rep-VLI), which achieves lightweight inference by reducing the computational overhead of text encoder.
  • We present the first fully spike–driven point Transformer. Its 3D spike–driven self–attention (3D-SDSA) reduces interactions to sparse additions, enabling faster, more efficient training that supports large–scale pretraining and generalizes broadly across 3D tasks.
@@ 2 Related Works
Numerous pretraining methods are proposed for spike-based representation learning. (Lee et al., 2018) utilized spike-timing-dependent plasticity (Bliss and Collingridge, 1993) to initialize SNNs, enhancing the model’s robustness and training speed. While this approach has been successful on simple datasets with shallow networks, its effectiveness diminishes as the complexity of the datasets and networks increases. To address these issues, SpikeBert and SpikeCLIP (Lv et al., 2023; Bal and Sengupta, 2024; Lv et al., 2025) employ a two-stage knowledge distillation process from ANNs to enhance spike-based representations for complex downstream tasks. However, these methods rely on ANN weight initialization, limiting structural flexibility. Additionally, they use LayerNorm, which hinders neuromorphic hardware deployment. SpikformerV2 and Spike-driven Transformer V3 (Yao et al., 2025; Zhou et al., 2024, 2023) apply a masked image modeling approach to address the performance degradation in SNNs as the model scales up. However, they require substantial storage and computational resources, and the lack of multimodal integration, particularly language guidance, limits their effectiveness in open-world understanding tasks.
@@
aim to align image and text embeddings for cross-modal transfer, with CLIP (Radford et al., 2021; Zhao et al., 2025; Chen et al., 2025; Li et al., 2021a) being a seminal work that uses contrastive learning for zero-shot classification. Building on this foundation, subsequent methods have expanded cross-modal alignment to include other modalities. These approaches typically fall into two categories: dual-encoder and triple-encoder frameworks. Dual-encoder fine-tune both visual and textual encoders (Lv et al., 2025; Zhang et al., 2021; Zhu et al., 2022). Triple-encoder frameworks incorporate additional modality-specific encoders (Xue et al., 2023, 2024; Zeng et al., 2023; Su et al., 2025), which combine triple models to achieve open-world understanding. This architecture is highly flexible, making it suitable for a variety of downstream tasks (Xu et al., 2024; Liu et al., 2023b; Wu et al., 2025a; Li et al., 2025). However, triple-encoder frameworks still rely on large text encoders during inference, hindering hardware deployment.
@@
from sparse, irregular data (events, point clouds) follows two main directions: point-based pipelines that operate on raw points to extract geometric features (Qi et al., 2017; Wang et al., 2019), and voxel-based pipelines that discretize into regular grids and apply sparse 3D convolutions (Wu et al., 2015). While deeper voxel/backbone designs can improve accuracy, the gains often come with substantial compute and memory costs, hindering deployment. To reduce cost, the SNN community has integrated spiking neurons with point-based models for low-power edge computing (Ren et al., 2024; Wu et al., 2024; Zhou et al., 2025); early designs, however, tend to be task-specific and capacity-limited. E-3DSNN (Qiu et al., 2025a) advances this line with spike sparse convolutions, delivering strong results across multiple 3D tasks while preserving spike-driven operation. Orthogonal to encoder choice, SVL pretraining enhances representation quality and enables open-world multimodal understanding while retaining spike efficiency.
Transformer-style spiking architectures are emerging. Spike Point Transformer (Wu et al., 2025b) introduces a Transformer-based SNN but still uses non-spiking operators and applies temporal encoding to point clouds, which degrades energy efficiency and slows training. Spike PointNet (Zhou et al., 2025) and E-3DSNN (Qiu et al., 2025a) rely on point-wise or sparse-convolutional inductive biases that may limit expressiveness and scalability. In contrast, we propose Spike PointFormer, the first fully spike-driven point Transformer: its 3D spike-driven self-attention (SDSA) performs addition-only interactions on spike tensors, enabling energy-efficient large-scale pretraining and broad generalization, and serving as a complementary architecture to SVL pretraining.
@@ 3 Preliminaries
are inspired by the dynamics of biological neurons (Maass, 1997; Li et al., 2023), which are the fundamental units of Spiking Neural Networks (SNNs). Among these, the Leaky Integrate-and-Fire (LIF) neuron is the most widely used due to its balance between biological plausibility and computational efficiency (Maass, 1997). We begin by translating the LIF spiking neuron into an iterative expression using the Euler method (Wu et al., 2018), which is described as follows:
@@
Directly training the above LIF-based SNNs requires the use of backpropagation through time (BPTT) (Wu et al., 2018), resulting in a time complexity of $\mathcal{O}(LT)$, where $L$ and $T$ are the number of layers and time steps. This significantly increases both the training time and memory requirements. To mitigate this issue, we use the Integer LIF Spiking Neuron.
@@
is incorporated into our SVL to reduce the quantization error, training time, and memory (Yao et al., 2025; Luo et al., 2024; Qiu et al., 2025b), which allows us to rewrite Eq. (2) as:
@@
where $\lfloor\cdot\rceil$ denotes the rounding operator, $\operatorname{clip}\{x,a,b\}$ confines $x$ within range $[a,b]$, and ${D}^{t}$ is a hyperparameter indicating the maximum emitted integer value by I-LIF. Moreover, I-LIF will emit integer values while pretraining and convert them into binary spikes by expanding the virtual timestep to ensure that the inference is spike-driven with only sparse addition.
@@ 4 Method
In this part, we present the 3D input representation, such as point clouds and event streams. Event streams, in particular, require special handling. We define them as $E_{i}=(x_{i},y_{i},t_{i},p_{i})$. Using a sliding window technique (Wang et al., 2019; Ren et al., 2024), we convert event streams into an event cloud, formulated as:
$E_{i}=(x_{i},y_{i},z_{i})\quad\text{where}\quad z_{i}=\frac{t_{i}-t_{\text{min}}}{t_{\text{max}}-t_{\text{min}}},$ (5)
By doing so, we treat event streams as a distinct kind of spatio-temporal point cloud. This allows us to consider both point clouds and event streams as collections of points, denoted by ${D}^{t}=\{\mathcal{P},\mathcal{F}\}$. This includes voxel sets ${D}_{k}^{t}=\{\mathcal{P}_{k}^{t},\mathcal{F}_{k}^{t}\}$, where $\mathcal{P}_{k}^{t}\in\mathbb{R}^{3}$ represents the 3D coordinates and $\mathcal{F}_{k}^{t}\in\mathbb{R}^{D}$ indicate the features across $d$ channels at the time step $t$. Following this, we utilize our I-LIF spiking neuron to encode these 3D inputs into spatio-temporal spike trains, which are then transmitted to the spike encoder.
@@ 4.2 Multi-scale Triple Alignment
To leverage the open-world recognition capabilities of the pre-trained CLIP model (Radford et al., 2021), we align the spike firing rate $\mathcal{F}^{S}/T$ with the text embeddings $\mathcal{F}^{T}$ obtained from CLIP, using a spike-text tuple $\mathcal{B}_{i}=\{{T}^{t}_{i},\mathcal{D}^{t}_{i}\}$ as input. The core idea is to bring the feature centroids of 3D instances and their corresponding text prompts closer together in the embedding space. To achieve this, we compute the InfoNCE loss (van den Oord et al., 2018) between the mean spike trains and the text features, as follows:
$\mathcal{L}^{\text{NCE}}_{(\textbf{S},\textbf{T})}=-\frac{1}{2|\mathcal{B}|}\sum_{i}^{|\mathcal{B}|}\log\frac{e^{\tau\mathbf{x}_{i}\cdot\mathbf{y}_{i}}}{\sum_{j}^{|\mathcal{B}|}e^{\tau\mathbf{x}_{i}\cdot\mathbf{y}_{j}}}+\log\frac{e^{\tau\mathbf{x}_{i}\cdot\mathbf{y}_{i}}}{\sum_{j}^{|\mathcal{B}|}e^{\tau\mathbf{x}_{j}\cdot y_{i}}},$ (6)
[…]
@@
$\mathcal{L}^{\text{NCE}}_{(\textbf{S},\textbf{I})}=-\frac{1}{2|\mathcal{C}|}\sum_{i}^{|\mathcal{C}|}\log\frac{e^{\tau\mathbf{a}_{i}\cdot\mathbf{b}_{i}}}{\sum_{j}^{|\mathcal{C}|}e^{\tau\mathbf{a}_{i}\cdot\mathbf{b}_{j}}}+\log\frac{e^{\tau\mathbf{a}_{i}\cdot\mathbf{b}_{i}}}{\sum_{j}^{|\mathcal{B}|}e^{\tau\mathbf{a}_{j}\cdot b_{i}}},$ (7)
[…]
@@
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are hyperparameters that balance the influence of image features, text features, and spike trains. For our main experiments, we set all three to 1, with a detailed ablation study on their impact presented in Tab 6.